40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 16 16:00:01 2019, gautam, Update, CDS, c1iscaux Supermicro setup IMG_7769.JPG
    Reply  Wed Jul 17 21:22:41 2019, gautam, Update, CDS, CM board Latch Enable subtlety 
       Reply  Sun Jul 21 12:55:38 2019, gautam, Update, CDS, CM board Latch Enable test script LatchLogic.pdfLatchLogicTest.png
          Reply  Fri Aug 9 08:59:04 2019, gautam, Update, CDS, Prep for install of c1iscaux 
             Reply  Sun Aug 11 11:47:42 2019, gautam, Update, CDS, Bench test of c1iscaux 
                Reply  Mon Aug 12 17:36:04 2019, gautam, Update, CDS, More bench test of c1iscaux iscauxCheclist.pdf
                   Reply  Mon Aug 12 21:25:19 2019, Koji, Update, CDS, More bench test of c1iscaux 
                   Reply  Tue Aug 13 08:07:09 2019, gautam, Update, CDS, P1--->P2 
                      Reply  Tue Aug 13 14:36:17 2019, gautam, Update, CDS, P1--->P2 
                   Reply  Fri Aug 16 16:40:04 2019, gautam, Update, CDS, 1Y3 work newLook1Y3.JPGIMG_7803.JPGc1lsc_crashed.png
                      Reply  Sat Aug 17 16:49:23 2019, gautam, Update, CDS, More 1Y3 work Screen_Shot_2019-08-17_at_3.00.57_PM.pngScreen_Shot_2019-08-17_at_3.12.23_PM.pngIMG_7804.JPGScreenshot_from_2019-08-17_17-04-47.png
                         Reply  Mon Aug 19 14:36:21 2019, gautam, Update, CDS, c1iscaux remaining work caseForSmallerFootprint.pdf
                            Reply  Tue Aug 20 19:05:24 2019, Koji, Update, CDS, MC1 (and MC3) troubleshoot Screenshot_from_2019-08-20_17-26-01.pngScreenshot_from_2019-08-20_17-43-03.png
                               Reply  Thu Aug 22 12:54:06 2019, Koji, Update, CDS, MC1 glitch removed (for now) and IMC locking recovered 7x
                                  Reply  Thu Aug 22 20:56:51 2019, Koji, Update, CDS, MC1 glitch removed (for now) and IMC locking recovered 
                            Reply  Fri Aug 23 18:46:17 2019, Jon, Update, CDS, c1iscaux remaining work 
                               Reply  Sun Aug 25 14:18:08 2019, gautam, Update, CDS, c1iscaux remaining work Screen_Shot_2019-08-25_at_10.38.37_PM.png
    Reply  Thu Jul 18 10:46:04 2019, gautam, Update, CDS, Database files made 
       Reply  Thu Jul 18 19:58:56 2019, gautam, Update, CDS, Work on Acromag chassis IMG_7771.JPGIMG_7770.JPGIMG_7772.JPG
          Reply  Fri Jul 19 19:44:03 2019, gautam, Update, CDS, Database file test Whitening.png
Message ID: 14849     Entry time: Sat Aug 17 16:49:23 2019     In reply to: 14848     Reply to this: 14850
Author: gautam 
Type: Update 
Category: CDS 
Subject: More 1Y3 work 

Work done today:

  1. All ribbon cable connections to the backplane of the 1Y2 Eurocrates were removed. The cables themselves were cleared for more space to work with.
  2. 20x 15ft DB37 Cables were run between 1Y2 and 1Y3 via overhead cable tray.
  3. Backplane interface boards were installed for 1Y2 Eurocrate boards.
  4. Connections were made between the Acromag chassis and the eurocrate electronics modules.

Testing of functionality:

  1. Fast BIO switching was verified to work for the following photodiodes:
    • AS55, AS110, REFL11, REFL33, REFL55, REFL165, POX11, POY11, POP22, POP110.
    • No light was incident on the PDs.
    • Test was done by increasing the whitening gain to +45 dB, and then looking at the ASD of the electronics noise between 50 Hz and 500 Hz with the whitening enabled/disabled. We expect x10 difference between the two states. This was seen.
  2. "DetMon" channels were verified to work - see Attachment #1
    • Y-axis units is volts
    • Test was done by toggling the output of the 11 MHz Marconi, and looking for a change.
    • As seen in the attachment, all 5 monitor channels show a change.
    • This needs to be calibrated into some sensible units - I don't know why the different modulation frequencies have such different readbacks from supposedly identical Demod Board monitor points.
    • Not sure if the ~10 V reported by the REFL165 monitor point is real or saturated.
    • These channels are installed to signal/help debug the infamous ERA-5 decay problem, but maybe already some are decayed?
  3. QPD interface channels were verified to work - see Attachment #2
    • Test was done by shining a green laser pointer on QPD quadrants.

Much testing remains to be done, but I defer further testing till Monday - the main functionality to be verified in the short run is the whitening gain stepping. The strain-relief of cables and general cleanup will be undertaken by Chub. Current state of affairs is in Attachment #3, leaves much to be desired in terms of cleanliness.

I will also setup the autoburt for the new machine on Monday. We will also need to add some channels to C0EDCU.ini if we want to trend them over some years (e.g. RF signal powers for monitoring ERA-5 health).

* c1lsc FE was rebooted using the usual script, and everything seems to be healthy in CDS-land again, see Attachment #4.

Quote:

Next steps: 

  1. We did not get around to running the DB37 cables between the Acromag chassis and the 1Y2 Eurocrates today - this operation itself took the whole day as we also needed to lay out some support struts etc on the rack to support the Sorensens and the Acromag chassis.
  2. Once the Acromags are connected to the Eurocrates, we have to run in-situ tests to make sure the appropriate functionality has been restored.
  3. We must have bumped something in the c1lsc expansion chassis - the CDS FE overview screen is reporting some errors (see Attachment #3). I will fix this.
  4. General tidiness, strain-relief etc.
Attachment 1: Screen_Shot_2019-08-17_at_3.00.57_PM.png  302 kB  Uploaded Sat Aug 17 20:27:12 2019  | Hide | Hide all
Screen_Shot_2019-08-17_at_3.00.57_PM.png
Attachment 2: Screen_Shot_2019-08-17_at_3.12.23_PM.png  140 kB  Uploaded Sat Aug 17 20:27:19 2019  | Hide | Hide all
Screen_Shot_2019-08-17_at_3.12.23_PM.png
Attachment 3: IMG_7804.JPG  1.562 MB  Uploaded Sat Aug 17 20:27:28 2019  | Hide | Hide all
IMG_7804.JPG
Attachment 4: Screenshot_from_2019-08-17_17-04-47.png  22 kB  Uploaded Sat Aug 17 20:27:34 2019  | Hide | Hide all
Screenshot_from_2019-08-17_17-04-47.png
ELOG V3.1.3-